OpenStudy (anonymous):

Show the process to evaluate the integral, int(x^3+x)/(x) dx, on the int, a 3 is on top and a 1 is on the bottom

OpenStudy (anonymous):

divide both terms by x : x^2 + 1 integrate each term seperately (1/3)x^3 + x substitute 3 for x and then 1 and subtract the results

OpenStudy (amistre64):

yeah, split the fraction into ist parts...

OpenStudy (amistre64):

\[\int\limits_{} \frac{x^3}{x}+ \frac{x}{x}dx\]

OpenStudy (anonymous):

okay, so would 4 be the answer

OpenStudy (amistre64):

from [1,3] [S] x^2 + 1 dx -> x^3/3 +x F(1) = 4/3 ; F(3) = 12 12 - 4/3 = 12 - 1'1/3 = 10'2/3

OpenStudy (amistre64):

I get 10' 2/3 ; or 32/3

OpenStudy (anonymous):

so no help...just checking